OpenStudy (anonymous):

Calculate the monthly finance charge for the following credit card transaction. Assume that it takes 10 days for a payment to be received and recorded and that the month is 30 days long. (Round your answer to the nearest cent.) $500 balance, 20% rate, $50 payment, previous balance method

OpenStudy (anonymous):

i do not know the formula to solve this can someone help

OpenStudy (anonymous):

can someone help

hero (hero):

@lalaly, help her

OpenStudy (lalaly):

PREVIOUS BALANCE METHOD: P=500 r=0.2 t=1/12 I=Prt =500*0.2*1/12

OpenStudy (anonymous):

thank you i couldnt figure out the method, i have several questions like these with no examples.

OpenStudy (lalaly):

I will try to help u if i can And ur welcome:) @Hero I help when i want to!

OpenStudy (anonymous):

Calculate the monthly finance charge for the following credit card transaction. Assume that it takes 10 days for a payment to be received and recorded and that the month is 30 days long. Assume 365 days in a year. (Round your answer to the nearest cent.) $3,000 balance, 21% rate, $1,500 payment, average daily balance method

hero (hero):

Yes, like now. You helped because you wanted to.

OpenStudy (anonymous):

would this be the same method

OpenStudy (lalaly):

yeah stay out of it @Hero

hero (hero):

@lalaly, you like being mean to me, don't you? :P

OpenStudy (lalaly):

@Hero no i dont .. Im sorry

OpenStudy (anonymous):

Thanks Shannon. So is the admjusted balance method 1500?

OpenStudy (anonymous):

can anyone answer this? Calculate the monthly finance charge for the credit card transaction. Assume that it takes 10 days for a payment to be received and recorded, and that the month is 30 days long. (Round your answers to the nearest cent.) $500 balance, 18%, $50 payment, previous balance method,adjusted balance method, average daily balance method